OpenStudy (anonymous):

Solve the inequality. 3m ≤ –9 (Points : 4) m ≥ –3 m ≤ –3 m ≤ –6 m ≥ 6

OpenStudy (anonymous):

You divide both sides by three so you get m ≤ -3.

OpenStudy (anonymous):

omg thank you much

OpenStudy (anonymous):

Welcome!

OpenStudy (anonymous):

can u help me with some more

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Solve the inequality. |x + 6| > 15 (Points : 4) x > –21 or x < 9 x < –21 or x > 9 –9 < x < 9 –21 < x < 9

OpenStudy (anonymous):

15 > x +6 > -15 9 > x > -21 soo the last one.

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

-21 < x < 9

OpenStudy (anonymous):

Solve the equation. |x + 5| = 9 (Points : 4) {–14, 4} {–4, 14} {4, 14} {–14, –4}

OpenStudy (anonymous):

x + 5 = 9 OR -(x +5) = 9 x = 4 or x = - 14

OpenStudy (anonymous):

So A.

OpenStudy (anonymous):

your so amzing thank you so much pretty and smart haha

OpenStudy (anonymous):

Hahah thanks. (: that is it?

OpenStudy (anonymous):

Which conjunction or disjunction is equivalent to the solution set? |x – 9| < 11 (Points : 4) x – 9 > 11 and x + 9 < –11 x – 9 < 11 and x – 9 > –11 x – 9 < 11 or x – 9 < –11 x – 9 < 11 or x – 9 > –11

OpenStudy (anonymous):

-11 < x -9 < 11 -2 < x < 20 x – 9 < 11 or x – 9 > –11 (D)

OpenStudy (anonymous):

Which is the solution set for –7 ≤ 4x + 1 ≤ 5? (Points : 4) x ≥ or x ≤ x ≥ and x ≤ x ≤ –2 or x ≥ 1 x ≥ –2 and x ≤ 1

OpenStudy (anonymous):

Those didn't show up..

OpenStudy (anonymous):

The length of a rectangle is twice its width. If the perimeter of the rectangle is less than 50 meters, what is the greatest width of the rectangle? (Points : 4) 2 • 2w + 2w < 50 2w • w < 50 2w < 50 2 + 2w < 50

OpenStudy (anonymous):

–7 ≤ 4x + 1 ≤ 5 -8 ≤ 4x ≤ 4 -2 ≤ x ≤ 1

OpenStudy (anonymous):

i dont need that one i need the one above u

OpenStudy (anonymous):

ples'

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

2 x 2w ?

OpenStudy (anonymous):

thats a

OpenStudy (anonymous):

2 • 2w + 2w < 50

OpenStudy (anonymous):

2. Solve and graph the inequality |4 – v| < 5. a. Write the inequality as two inequalities without absolute value. b. Solve the inequality and write the solution set. c. Graph the solution on a number line. 2. Solve and graph the inequality |4 – v| < 5. a. Write the inequality as two inequalities without absolute value. b. Solve the inequality and write the solution set. c. Graph the solution on a number line. 2. Solve and graph the inequality |4 – v| < 5.